By a suitable change of origin, show that the equation 9(x − 5)2 + 16(y + 3)2 = 144 represent an ellipse and that the coordinates of the centre are (5, −3). Find the position vectors of the foci and the ends of the major and minor axes.


1
Expert's answer
2021-08-15T17:35:15-0400

"9(x-5)^2+16(y+3)^2=144"


"\dfrac{(x-5)^2}{16}+\dfrac{(y+3)^2}{9}=1"

The equation of ellipse


"\dfrac{(x-a_C)^2}{a^2}+\dfrac{(y-y_C)^2}{b^2}=1"

The center "C(5, -3)"


"a=4, b=3"


"c=\sqrt{a^2-b^2}=\sqrt{4^2-3^2}=\sqrt{7}"

Foci: "(5-\sqrt{7}, -3), (5+\sqrt{7}, -3)."


Ends of the major axis: "(1, -3), (9, -3)."


Ends of the minor axis: "(5, -6), (5, 0)."